# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

### v1

@@ G.2 Corpora
  • RealNews/Regional is a subset of RealNews (Zellers et al., 2019). It includes 2.7M articles from non-US and non-EU websites. We manually check RealNews websites and identified 141 regional-news websites with headquarters in 56 non-US and non-EU countries: India, Pakistan, Nigeria, Philippines, etc.
  • RealNews/India is further filtered to only include Indian news websites.
@@
  • Products: We pull the data from https://nijianmo.github.io/amazon/index.html#complete-data and concatenate title and description
Appendix H Teacher and Student hyperparameters
H.1 Teacher LLM hyperparams
@@ H.2 Student LM hyperparams
H.3 Oracle model hyperparams
@@ H.4 Retriever
Appendix I Computational budget
We run all our models on AWS Elastic Cloud Compute [^2] https://aws.amazon.com/ec2/ using 20 p3dn.24xlarge machines.
I.1 Information Retrieval
@@
I.2 Dataset synthesis
For LLaMa-2 Chat 13B, we create 48 copies of the model and orchestrate it using the Ray distributed framework. Generation is done in roughly 6 hours per dataset of 8k rows.
To use Claude Instant-v1, we invoke AWS Bedrock [^5] https://docs.aws.amazon.com/pdfs/bedrock/latest/APIReference/bedrock-api.pdf using the boto3 library [^6] https://boto3.amazonaws.com/v1/documentation/api/latest/index.html. Generation from Claude was done at an AWS-account level RPM of 1600 and takes roughly 4 hours per dataset on 8k rows.
I.3 Student distillation
Each DeBERTa-v3-Large student model trains for between 1 to 3 hours on a single V100 GPU on 8k rows. Each DistilBERT student model trains in 1 hour to generate the data-map.
Appendix J Licensing

### v3

@@ I.2 Corpora
  • RealNews/Regional is a subset of RealNews (Zellers et al., 2019). It includes 2.7M articles from non-US and non-EU websites. We manually check RealNews websites and identified 141 regional-news websites with headquarters in 56 non-US and non-EU countries: India, Pakistan, Nigeria, Philippines, etc. The complete list is mentioned in Table 27.
  • RealNews/India is further filtered to only include Indian news websites. We use only the “India” domains from Table 27.
@@
  • Products: We pull the data from https://nijianmo.github.io/amazon/index.html#complete-data and concatenate title and description.
  • CMU Movie Summary: Data is obtained from https://www.cs.cmu.edu/~ark/personas/, where we use the plot summaries file.
Appendix J Teacher and Student hyperparameters
J.1 Teacher LLM hyperparams
@@
J.3 Oracle model hyperparams
@@
Appendix K Computational budget
We run all our models on AWS Elastic Cloud Compute [^3] https://aws.amazon.com/ec2/ using 20 p3dn.24xlarge machines to call AWS cloud services, host the retrieval index and distill student models.
K.1 Information Retrieval
@@
K.2 Dataset synthesis
In order to run LLaMa-2 Chat 13B and Claude Instant-v1, we invoke AWS Bedrock [^6] https://docs.aws.amazon.com/pdfs/bedrock/latest/APIReference/bedrock-api.pdf using the boto3 library [^7] https://boto3.amazonaws.com/v1/documentation/api/latest/index.html.
Generations were done at an AWS-account level RPM of 1600 and takes roughly 4 hours for a dataset of 8k rows.
K.3 Student distillation
Each DeBERTa-v3-Large student model trains for 1-3 hours on a single GPU on 8k rows. Each DistilBERT student model trains in 1 hour to generate the data-map for dataset catrography.
Appendix L Licensing